if a number is divisible by 40 and remainder is 30 what is the remainder is the number is divisible by 20
Answer by richwmiller(17219)

since any number divisible by 40 is also divisible by 20 the answer would be double. Plus 30 has one 20 in it.
So the answer would be double the original plus 1 with a remainder of 10.
30/20=1 plus remainder of 10
let's take 150
150/40=3 with a remainder of 30
150/20=6 plus 1 which is 7 with a remainder of 10.